OpenStudy (anonymous):

Finding maximum and minimum values for the function for the polygonal convex

OpenStudy (anonymous):

http://prntscr.com/8h1ig1

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

I tried one point and I got this f(2,0)=4(2)+7(0)=8+0=8